OpenStudy (12wishs):

Can some one help!! Will Medal! Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)=x+ab g(x)=cx−d

OpenStudy (12wishs):

please help

OpenStudy (alekos):

Tough one

OpenStudy (12wishs):

ya

OpenStudy (alekos):

Have they given you an answer?

OpenStudy (12wishs):

no I have to imput my own numbers into the functions given and figure it out my self and this is only part one

OpenStudy (12wishs):

umm.... you still there?

OpenStudy (alekos):

Yeah. Trying to work it out algebraically

OpenStudy (12wishs):

ok

OpenStudy (alekos):

The inverse of y=x + ab is y = x - ab So for g(x) = cx - d we get c=1 and ab=d

OpenStudy (alekos):

a and b can be any real number

OpenStudy (12wishs):

ok thx